# Taming the Long-Tail: Efficient Reasoning RL Training with Adaptive Drafter

### v1

@@ 1. Introduction
Over the past year, we have witnessed the emergence of reasoning LLMs, such as OpenAI-o1 (O1), DeepSeek-R1 (DeepSeekR1) and Gemini-2.5Pro (Gemini2.5Pro), exhibiting remarkable capability improvements in mathematical reasoning, programming, and multidisciplinary knowledge tasks (ReasoningSurvey). A key driver behind these improvements is the use of reinforcement learning (RL), which has become the standard approach for endowing LLMs with strong reasoning capabilities (DeepSeekMath; QwQ; DeepSeekR1; DAPO). However, this RL approach faces efficiency challenges due to the unique workload characteristics involved. Our analysis of self-collected traces (Figure 1) and production traces from ByteDance (Figure 2) identifies the following key characteristics of reasoning RL process:
@@
(2) Persistent Long-Tail Distribution. A major source of RL training inefficiency arises from the long-tail distribution of rollout response lengths. As shown in Figure 1(a), although most generated sequences are relatively short, a small fraction extends to extreme lengths. Crucially, this inefficiency is not an occasional occurrence but a persistent pattern observed throughout long-term training, as evidenced by ByteDance’s production traces (Figure 2). In most RL steps, a few responses reach the maximum configured length, while the majority remain substantially shorter. The pronounced gap between this maximum length and the 75th percentile (p75) underscores significant resource under-utilization.
@@
Existing systems (HybridFlow; RLHFuse; ReaLHF; DeepSpeedChat; GEAR; NeMoAligner; FlexRLHF) for RLHF (RL from Human Feedback) primarily focus their optimization efforts on managing multiple models, optimizing data transfer, and orchestrating device allocation across the different phases. However, they neglect tackling the rollout bottleneck, thereby remaining inefficient for reasoning RL tasks. Critically, reasoning rollouts are often over an order of magnitude longer than typical RLHF outputs (HybridFlow; RLHFuse; InstructGPT), giving reasoning RL fundamentally different workload characteristics and demanding distinct system-level optimizations.
@@
While various acceleration techniques exist for LLM serving, many are ill-suited for the RL training context. Methods like quantization (AWQ; GPTQ) and model sparsity (LLM-Pruner; StreamingLLM), despite offering speedups, are typically lossy, risking model accuracy degradation and altering output probabilities. We identify Speculative Decoding (SD) (SpeculativeDecoding; chen2023accelerating) as a highly suitable approach. SD utilizes a lightweight draft model to generate token sequences speculatively, which are then verified in parallel by the main target model. SD is particularly well-suited for RL training due to two primary merits: (1) Mathematically Lossless. The output distribution remains identical to the target model’s original distribution; (2) Efficient for Long-Tail. SD enhances throughput by shifting the process from being memory-bound towards compute-bound, which is particularly effective for the long-tail phase of RL rollouts, where effective batch sizes are typically small.
@@
However, existing speculative decoding techniques (Medusa; Eagle1; SpecInfer; HASS; Falcon; REST) are designed for static inference scenarios, where the target model is fixed. Applying them within the dynamic reasoning RL training is fundamentally non-trivial, presenting significant challenges: (1) Evolving Target Model: the target model’s weights are constantly updated during RL training, making the draft model progressively stale, severely undermining the effectiveness of SD. (2) Draft Model Training Costs: Effective SD often relies on specialized draft models (HASS; Eagle1; Eagle2; Eagle3; Falcon; Hydra), which require extra training to align with the target model, introducing additional overhead and complexity for users. (3) Scheduling Complexity: SD performance is sensitive to batch size and hyperparameters (e.g., draft depth). Reasoning RL rollouts, however, are characterized by dynamically fluctuating effective batch sizes as sequences complete at different times, necessitating an effective strategy selector for efficient SD deployment.
@@
To address these challenges, we introduce TLT, an efficient system for reasoning RL training featuring adaptive speculative decoding. TLT aims to mitigate the long-tail rollout bottleneck while guaranteeing mathematical losslessness. The core design of TLT derives from the following insights: First, Exploiting Rollout Bubbles. The characteristically long duration of RL rollouts provides ample time to utilize GPU resources that progressively become available due to the long-tail nature. Resources freed as sequences complete can thus be repurposed for other tasks, such as draft model training, without additional costs. Second, Non-Blocking Drafter Training. Draft model training can be decoupled from full rollout completion; training asynchronously on partial or readily available data allows its computation to be effectively overlapped with ongoing rollouts.
@@ 2.1. Reasoning and Reinforcement Learning
Reasoning LLMs. Large Language Models (LLMs) have recently shown remarkable performance in complex reasoning domains such as mathematics, logic, and programming. Powerful models such as OpenAI-o1 (O1) and DeepSeek-R1 (DeepSeekR1) achieve impressive accuracy on challenging reasoning benchmarks, including competition-level math and coding challenges. A key factor in their success is test-time scaling (InferenceScalingLaws; GreedyPolicySearch), an inference paradigm that enhances accuracy by affording models more generation time to explore complex reasoning paths. These reasoning LLMs typically feature: (1) Long Chain-of-Thought, generating extended intermediate reasoning steps before reaching a final answer; (2) Self-Reflection, the capacity to evaluate, refine, and correct their own reasoning during inference to improve accuracy. As illustrated in Figure 3 (a), allowing extended generation time enables these models to explore deeper and wider thinking traces (ReasoningSurvey), thereby enhancing accuracy without additional training costs. More importantly, unlike traditional LLMs, these reasoning models can automatically revisit their generated responses and exhibit self-reflection capabilities (e.g., ”wait” in Figure 3 (b) reasoning trace), which are typically acquired during their RL training stage (DeepSeekR1) and are key to performance enhancements.
Reinforcement Learning for Reasoning. Recently, RL algorithm variants such as Group Relative Policy Optimization (GRPO) (DeepSeekMath) have proven their success in training reasoning LLMs. Unlike standard Policy Optimization (PPO) (PPO), which relies on an extensive critic model, GRPO employs a group-based baseline estimation to streamline the optimization process, significantly reducing training overhead while preserving the robustness of policy updates (PostTrainingSurvey).
Figure 4 illustrates a single training step in GRPO, which comprises three main stages: (1) Rollout Stage, the target model generates multiple candidate responses for each input prompt; (2) Inference Stage, the generated responses are processed by both the target model (the model undergoing RL training) and a fixed reference model (the initial target model, i.e., RL step=0). This processing yields logits used to compute a KL divergence penalty (john2020kld), which serves to constrain model updates, ensuring stability and preventing excessive deviation from the reference model. Concurrently within this stage, a reward is calculated with a rule-based policy regarding response characteristics (e.g., correctness), rather than relying on a separate value model; (3) Training Stage, a loss function is constructed using the reward value and KL divergence. The target model’s weights are then updated according to this loss.
@@ 2.2. Speculative Decoding for Efficient Reasoning RL
As mentioned in §1, the long-tail rollout stage in reasoning RL frequently becomes a performance bottleneck. Consequently, tailored system optimizations for decoding acceleration are essential to improve throughput and resource efficiency. To this end, we identify speculative decoding (SpeculativeDecoding; chen2023accelerating) as a key approach well-suited to tackle these challenges.
@@
C2. Draft Model Training Costs. Achieving high SD efficiency often requires dedicated draft models (e.g., single-decoder-layer models like Eagle (Eagle1), HASS (HASS)), rather than arbitrary smaller models. Training a specialized drafter for the target model introduces substantial additional cost.
@@ 3. TLT Overview
(d) Generalizable and Scalable. The design must be adaptable to various popular reasoning RL algorithms (e.g., GRPO (DeepSeekMath), RLOO (RLOO)). Furthermore, TLT should scale effectively and deliver continuous system performance improvements across different model sizes, architectures, and cluster scales.
@@ 4. Adaptive Drafter
| Strategies | ④HiddenStates | ⑤LossFunc |
| Eagle (Eagle1) | Last layer | L1 + CE loss | N.A. | ||
| HASS (HASS) | Last layer | L1 + CE loss | 3 steps | ||
| Eagle-3 (Eagle3) |
CE loss only | 7 steps |
@@ 4.1. Draft Model
Model Architecture. To maximize efficiency, TLT adopts a highly lightweight, single-layer model design for the drafter (Eagle1; HASS; Eagle3). In this context, a ”layer” refers to one complete transformer decoder block, including attention and feed-forward components. This approach differs from vanilla speculative decoding (SpeculativeDecoding; chen2023accelerating), which often uses a separate, smaller multi-layer LLM from the same family (e.g., Qwen2.5-0.5B for a Qwen2.5-32B target). The vanilla approach has drawbacks: suitable and accurate small drafters may not exist for every target LLM, not to mention that these smaller models can still incur substantial drafting latency. For instance, while Qwen2.5-32B has 64 layers, Qwen2.5-0.5B still contains 24 layers. Although the parameter count is reduced significantly (64 in this case), latency remains dominated by the sequential computation across multiple layers. Experiments demonstrate that TLT’s single-layer draft model, sharing approximately the same parameter count as the Qwen2.5-0.5B drafter, is significantly faster (2.4).
As illustrated in Figure 7 (blue block), the TLT drafter mirrors the target model’s architecture but incorporates only a single, trainable decoder layer. It reuses the target model’s original Embedding and LM Head layers, keeping their weights frozen during training. Consequently, only the single decoder layer’s parameters are updated, representing a small fraction (approximately 1/layer_num) of the target model’s total parameters. This architectural design significantly reduces both training and inference overhead. The effectiveness of such single-layer drafters is corroborated by prior work (HASS; Eagle1; Eagle2; Eagle3; Falcon; Hydra), which has shown their capability to closely align with the target model’s auto-regressive distribution and achieve high acceptance rates for speculative tokens. During operation, the trained drafter takes hidden states from the target model and auto-regressively drafts subsequent tokens for verification.
@@
Additionally, several training strategies are supported within this unified pipeline in TLT. For instance, EAGLE (Eagle1) uses only the last layer’s hidden states; HASS (HASS) further enhance the training by feeding the output feature of draft model back into itself for multiple time (i.e., training-time test); EAGLE-3 (Eagle3) further fuses multi-layer hidden states from the target model, aiming for better acceptance rates. Despite these differences, all these variants can be naturally expressed within our training framework. In this work, we use the EAGLE model as the default, as it offers high accepted lengths and faster convergence with much lower training cost.
@@ 4.2. Spot Trainer
Worker Coordinator. It aims to bridge the gap between the Rollout Engine and the Spot Trainer, coordinating resource allocation, particularly repurposing GPU resources that become idle during the long-tail phase of rollout. In TLT, a worker is the basic unit of resource management, defined as one rollout instance. For example, when rollout uses a full DGX-H100 node with TP=8, a single worker represents 8 GPUs. The coordinator follows a centralized model implemented with ZeroMQ (ZeroMQ), where a single coordinator process (rank 0) maintains the global state and workers communicate through asynchronous request-reply patterns. By monitoring active request counts and worker states, the coordinator determines when workers can be migrated from rollout to opportunistic drafter training. Each worker cycles between three states—BUSY (serving rollout), IDLE (inference finished and memory released), and TRAINING (engaged in drafter updates)—and notifies the coordinator on every transition. Once the number of idle workers exceeds a configurable threshold, the coordinator promotes them to training mode and initiates drafter training.
@@
Selective Asynchronous Checkpointing. A key design feature of the Spot Trainer is its preemptibility, allowing it to utilize idle resources without blocking the main RL workload. However, preemptions may lead to substantial loss of the draft model’s training progress. To address this, the Spot Trainer utilizes asynchronous checkpointing: when triggered, the process of saving the draft model’s state is offloaded to a background thread. Moreover, we optimize this process by filtering out frozen layers (e.g., embeddings and the LM head) and dumping only the trainable parameters, significantly reducing checkpointing latency. This mechanism minimizes work lost due to preemption through frequent and efficient checkpointing.
Sequence Packing. The training data inherently consists of sequences with variable lengths. Standard batching techniques often require padding sequences to a uniform length, leading to inefficiency of computation, communication, and memory. To maximize GPU utilization during opportunistic drafter training, the Spot Trainer employs sequence packing (SeqPacking). This technique concatenates multiple variable-length sequences into a single packed sequence, removing padding while using attention masks to preserve sequence integrity during parallel processing. This enables efficient processing of the unbalanced training data within the limited, preemptible time slots available for spot training.
@@ 5.1. Adaptive SD Manager
Tree-based Drafting. Unlike linear drafting, which processes a single sequence, tree-based drafting extends this single-branch paradigm by enabling parallel exploration of multiple speculative paths. Following (SpecInfer; Eagle2), we build the candidate tree based on the draft model’s confidence score (token probability), as illustrated in Figure 9. Starting from the input prompt (”Climate”), the draft model predicts potential next tokens. The tree-based approach explores the topK most likely options at each step (e.g., topK=2 yields ”change” and ”control”), and this branching continues for Draft_Depth steps, creating a tree of candidate sequences. Finally, the Tokens_to_Verify=8 highest-confidence tokens within the tree are chosen and submitted to the target model for parallel verification. By exploring and verifying multiple potential sequences simultaneously, tree-based drafting substantially increases the number of accepted tokens per verification step, thereby improving overall generation throughput with more effective utilization of parallel compute resources.
Adaptivity is Necessary. While the SD is primarily designed to accelerate the long-tail stage, we observe that SD can also deliver speedups at moderate batch sizes (Table 2). However, this requires adaptive adjustment of SD strategies according to the decoding batch size and workload characteristics. Prior work (Eagle1; Eagle2; HASS) mainly focuses on batchsize=1 scenarios, which limits its applicability in real-world scenarios. RL rollout naturally involve highly dynamic batch sizes: they begin large and then shrink rapidly as short responses complete. Static SD strategies are therefore both inefficient and unsafe, as they may cause Out-of-Memory (OOM) failures under large batches. To address this, the Adaptive SD Manager continuously tunes SD parameters—choosing safe strategies to prevent OOM, maximizing throughput by learning effective strategies for each workload, and balancing speculative gains against computational overhead. In addition, to handle large batch sizes where SD offers little benefit, it employs an elastic mechanism that activates SD only when the remaining request count drops below a configurable threshold (default: 32).

### v2

@@ 1. Introduction
Over the past year, we have witnessed the emergence of reasoning LLMs, such as OpenAI-o1 (OpenAI et al., 2025), DeepSeek-R1 (DeepSeek-AI et al., 2025) and Gemini-2.5Pro (Gemini, 2025), exhibiting remarkable capability improvements in mathematical reasoning, programming, and multidisciplinary knowledge tasks (Chen et al., 2025). A key driver behind these improvements is the use of reinforcement learning (RL), which has become the standard approach for endowing LLMs with strong reasoning capabilities (Shao et al., 2024; Team, 2025; DeepSeek-AI et al., 2025; Yu et al., 2025). However, this RL approach faces efficiency challenges due to the unique workload characteristics involved. Our analysis of self-collected traces (Figure 1) and production traces from ByteDance (Figure 2) identifies the following key characteristics of reasoning RL process:
@@
(2) Persistent Long-Tail Distribution. A major source of RL training inefficiency arises from the long-tail distribution of rollout response lengths. As shown in Figure 1(a), although most generated sequences are relatively short, a small fraction extends to extreme lengths. Crucially, this inefficiency is not an occasional occurrence but a persistent pattern observed throughout long-term training, as evidenced by ByteDance’s production traces (Figure 2). In most RL steps, a few responses reach the maximum configured length, while the majority remain substantially shorter. The pronounced gap between this maximum length and the 75th percentile (p75) underscores significant resource under-utilization. Similar long-tail patterns have been independently reported in large-scale RL systems at Berkeley (Luo et al., 2025), Alibaba (Gao et al., 2025a), and AMD (Zhou et al., 2025), indicating that this phenomenon is widespread.
@@
Existing systems (Sheng et al., 2025; Zhong et al., 2025b; Mei et al., 2024; Yao et al., 2023; Wang et al., 2023; Shen et al., 2024; Xiao et al., 2024b) for RLHF (RL from Human Feedback) primarily focus their optimization efforts on managing multiple models, optimizing data transfer, and orchestrating device allocation across the different phases. However, they neglect tackling the rollout bottleneck, thereby remaining inefficient for reasoning RL tasks. Critically, reasoning rollouts are often over an order of magnitude longer than typical RLHF outputs (Sheng et al., 2025; Zhong et al., 2025b; Ouyang et al., 2022), giving reasoning RL fundamentally different workload characteristics and demanding distinct system-level optimizations.
@@
While various acceleration techniques exist for LLM serving, many are ill-suited for the RL training context. Methods like quantization (Lin et al., 2024; Frantar et al., 2023) and model sparsity (Ma et al., 2023; Xiao et al., 2024a), despite offering speedups, are typically lossy, risking model accuracy degradation and altering output probabilities. We identify Speculative Decoding (SD) (Leviathan et al., 2023; Chen et al., 2023) as a highly suitable approach. SD utilizes a lightweight draft model to generate token sequences speculatively, which are then verified in parallel by the main target model. SD is particularly well-suited for RL training due to two primary merits: (1) Mathematically Lossless. The output distribution remains identical to the target model’s original distribution; (2) Efficient for Long-Tail. SD enhances throughput by shifting the process from being memory-bound towards compute-bound, which is particularly effective for the long-tail phase of RL rollouts, where effective batch sizes are typically small.
@@
However, existing speculative decoding techniques (Cai et al., 2024; Li et al., 2024b; Miao et al., 2024; Zhang et al., 2025; Gao et al., 2025b; He et al., 2024) are designed for static inference scenarios, where the target model is fixed. Applying them within the dynamic reasoning RL training is fundamentally non-trivial, presenting significant challenges: (1) Evolving Target Model: the target model’s weights are constantly updated during RL training, making the draft model progressively stale, severely undermining the effectiveness of SD. (2) Draft Model Training Costs: Effective SD often relies on specialized draft models (Zhang et al., 2025; Li et al., 2024b, a, 2025a; Gao et al., 2025b; Ankner et al., 2024), which require extra training to align with the target model, introducing additional overhead and complexity for users. (3) Scheduling Complexity: SD performance is sensitive to batch size and hyperparameters (e.g., draft depth). Reasoning RL rollouts, however, are characterized by dynamically fluctuating effective batch sizes as sequences complete at different times, necessitating an effective strategy selector for efficient SD deployment.
@@
To address these challenges, we introduce TLT, an efficient system for reasoning RL training featuring adaptive speculative decoding. TLT aims to mitigate the long-tail rollout bottleneck while guaranteeing mathematical losslessness. The core design of TLT derives from the following insights: First, Exploiting Rollout Bubbles. The characteristically long duration of RL rollouts provides ample time to utilize GPU resources that progressively become available due to the long-tail nature. Resources freed as sequences complete can thus be repurposed for other tasks, such as draft model training, without additional costs (Hu et al., 2023). Second, Non-Blocking Drafter Training. Draft model training can be decoupled from full rollout completion; training asynchronously on partial or readily available data allows its computation to be effectively overlapped with ongoing rollouts.
@@ 2.1. Reasoning and Reinforcement Learning
Reasoning LLMs. Large Language Models (LLMs) have recently shown remarkable performance in complex reasoning domains such as mathematics, logic, and programming. Powerful models such as OpenAI-o1 (OpenAI et al., 2025) and DeepSeek-R1 (DeepSeek-AI et al., 2025) achieve impressive accuracy on challenging reasoning benchmarks, including competition-level math and coding challenges. A key factor in their success is test-time scaling (Wu et al., 2025; Lyzhov et al., 2020), an inference paradigm that enhances accuracy by affording models more generation time to explore complex reasoning paths. These reasoning LLMs typically feature: (1) Long Chain-of-Thought, generating extended intermediate reasoning steps before reaching a final answer; (2) Self-Reflection, the capacity to evaluate, refine, and correct their own reasoning during inference to improve accuracy. As illustrated in Figure 3 (a), allowing extended generation time enables these models to explore deeper and wider thinking traces (Chen et al., 2025), thereby enhancing accuracy without additional training costs. More importantly, unlike traditional LLMs, these reasoning models can automatically revisit their generated responses and exhibit self-reflection capabilities (e.g., ”wait” in Figure 3 (b) reasoning trace), which are typically acquired during their RL training stage (DeepSeek-AI et al., 2025) and are key to performance enhancements.
Reinforcement Learning for Reasoning. Recently, RL algorithm variants such as Group Relative Policy Optimization (GRPO) (Shao et al., 2024) have proven their success in training reasoning LLMs. Unlike standard Policy Optimization (PPO) (Schulman et al., 2017), which relies on an extensive critic model, GRPO employs a group-based baseline estimation to streamline the optimization process, significantly reducing training overhead while preserving the robustness of policy updates (Tie et al., 2025).
Figure 4 illustrates a single training step in GRPO, which comprises three main stages: (1) Rollout Stage, the target model generates multiple candidate responses for each input prompt; (2) Inference Stage, the generated responses are processed by both the target model (the model undergoing RL training) and a fixed reference model (the initial target model, i.e., RL step=0). This processing yields logits used to compute a KL divergence penalty (Schulman, 2020), which serves to constrain model updates, ensuring stability and preventing excessive deviation from the reference model. Concurrently within this stage, a reward is calculated with a rule-based policy regarding response characteristics (e.g., correctness), rather than relying on a separate value model; (3) Training Stage, a loss function is constructed using the reward value and KL divergence. The target model’s weights are then updated according to this loss.
@@ 2.2. Speculative Decoding for Efficient Reasoning RL
As mentioned in §1, the long-tail rollout stage in reasoning RL frequently becomes a performance bottleneck. Consequently, tailored system optimizations for decoding acceleration are essential to improve throughput and resource efficiency. To this end, we identify speculative decoding (Leviathan et al., 2023; Chen et al., 2023) as a key approach well-suited to tackle these challenges.
@@
C2. Draft Model Training Costs. Achieving high SD efficiency often requires dedicated draft models (e.g., single-decoder-layer models like Eagle (Li et al., 2024b), HASS (Zhang et al., 2025)), rather than arbitrary smaller models. Training a specialized drafter for the target model introduces substantial additional cost.
@@ 3. TLT Overview
(d) Generalizable and Scalable. The design must be adaptable to various popular reasoning RL algorithms (e.g., GRPO (Shao et al., 2024), RLOO (Ahmadian et al., 2024)). Furthermore, TLT should scale effectively and deliver continuous system performance improvements across different model sizes, architectures, and cluster scales.
@@ 4.1. Draft Model
Model Architecture. To maximize efficiency, TLT adopts a highly lightweight, single-layer model design for the drafter (Li et al., 2024b; Zhang et al., 2025; Li et al., 2025a). In this context, a ”layer” refers to one complete transformer decoder block, including attention and feed-forward components. This approach differs from vanilla speculative decoding (Leviathan et al., 2023; Chen et al., 2023), which often uses a separate, smaller multi-layer LLM from the same family (e.g., Qwen2.5-0.5B for a Qwen2.5-32B target). The vanilla approach has drawbacks: suitable and accurate small drafters may not exist for every target LLM, not to mention that these smaller models can still incur substantial drafting latency. For instance, while Qwen2.5-32B has 64 layers, Qwen2.5-0.5B still contains 24 layers. Although the parameter count is reduced significantly (64 in this case), latency remains dominated by the sequential computation across multiple layers. Experiments demonstrate that TLT’s single-layer draft model, sharing approximately the same parameter count as the Qwen2.5-0.5B drafter, is significantly faster (2.4).
As illustrated in Figure 7 (blue block), the TLT drafter mirrors the target model’s architecture but incorporates only a single, trainable decoder layer. It reuses the target model’s original Embedding and LM Head layers, keeping their weights frozen during training. Consequently, only the single decoder layer’s parameters are updated, representing a small fraction (approximately 1/layer_num) of the target model’s total parameters. This architectural design significantly reduces both training and inference overhead. The effectiveness of such single-layer drafters is corroborated by prior work (Zhang et al., 2025; Li et al., 2024b, a, 2025a; Gao et al., 2025b; Ankner et al., 2024), which has shown their capability to closely align with the target model’s auto-regressive distribution and achieve high acceptance rates for speculative tokens. During operation, the trained drafter takes hidden states from the target model and auto-regressively drafts subsequent tokens for verification.
@@
Additionally, several training strategies are supported within this unified pipeline in TLT, as our framework is drafter-agnostic. For instance, EAGLE (Li et al., 2024b) uses only the last layer’s hidden states; HASS (Zhang et al., 2025) further enhance the training by feeding the output feature of draft model back into itself for multiple time (i.e., training-time test); EAGLE-3 (Li et al., 2025a) further fuses multi-layer hidden states from the target model, aiming for better acceptance rates. Despite these differences, all these variants can be naturally expressed within our training framework. In this work, we use the EAGLE model as the default, as it offers high accepted lengths and faster convergence with much lower training cost.
@@ 4.2. Spot Trainer
Worker Coordinator. It aims to bridge the gap between the Rollout Engine and the Spot Trainer, coordinating resource allocation, particularly repurposing GPU resources that become idle during the long-tail phase of rollout. In TLT, a worker is the basic unit of resource management, defined as one rollout instance. For example, when rollout uses a full DGX-H100 node with TP=8, a single worker represents 8 GPUs. The coordinator follows a centralized model implemented with ZeroMQ (Hintjens, 2013), where a single coordinator process (rank 0) maintains the global state and workers communicate through asynchronous request-reply patterns. By monitoring active request counts and worker states, the coordinator determines when workers can be migrated from rollout to opportunistic drafter training. Each worker cycles between three states—BUSY (serving rollout), IDLE (inference finished and memory released), and TRAINING (engaged in drafter updates)—and notifies the coordinator on every transition. Once the number of idle workers exceeds a configurable threshold, the coordinator promotes them to training mode and initiates drafter training.
@@
Selective Asynchronous Checkpointing. A key design feature of the Spot Trainer is its preemptibility, allowing it to utilize idle resources without blocking the main RL workload. However, preemptions may lead to substantial loss of the draft model’s training progress. To address this, the Spot Trainer utilizes asynchronous checkpointing: when triggered, the process of saving the draft model’s state is offloaded to a background thread. Moreover, we optimize this process by filtering out frozen layers (e.g., embeddings and the LM head) and dumping only the trainable parameters, significantly reducing checkpointing latency. This mechanism minimizes work lost due to preemption through frequent and efficient checkpointing (Hu et al., 2024).
Sequence Packing. The training data inherently consists of sequences with variable lengths. Standard batching techniques often require padding sequences to a uniform length, leading to inefficiency of computation, communication, and memory. To maximize GPU utilization during opportunistic drafter training, the Spot Trainer employs sequence packing (Krell et al., 2022). This technique concatenates multiple variable-length sequences into a single packed sequence, removing padding while using attention masks to preserve sequence integrity during parallel processing. This enables efficient processing of the unbalanced training data within the limited, preemptible time slots available for spot training.
@@ 5.1. Adaptive SD Manager
Tree-based Drafting. Unlike linear drafting, which processes a single sequence, tree-based drafting extends this single-branch paradigm by enabling parallel exploration of multiple speculative paths. Following (Miao et al., 2024; Li et al., 2024a), we build the candidate tree based on the draft model’s confidence score (token probability), as illustrated in Figure 9. Starting from the input prompt (”Climate”), the draft model predicts potential next tokens. The tree-based approach explores the topK most likely options at each step (e.g., topK=2 yields ”change” and ”control”), and this branching continues for Draft_Depth steps, creating a tree of candidate sequences. Finally, the Tokens_to_Verify=8 highest-confidence tokens within the tree are chosen and submitted to the target model for parallel verification. By exploring and verifying multiple potential sequences simultaneously, tree-based drafting substantially increases the number of accepted tokens per verification step, thereby improving overall generation throughput and GPU utilization.
Adaptivity is Necessary. While the SD is primarily designed to accelerate the long-tail stage, we observe that SD can also deliver speedups at moderate batch sizes (Table 4). However, this requires adaptive adjustment of SD strategies according to the decoding batch size and workload characteristics. Prior work (Li et al., 2024b, a; Zhang et al., 2025) mainly focuses on batchsize=1 scenarios, which limits its applicability in real-world scenarios. RL rollout naturally involve highly dynamic batch sizes: they begin large and then shrink rapidly as short responses complete. Static SD strategies are therefore both inefficient and unsafe, as they may cause Out-of-Memory (OOM) failures under large batches. To address this, the Adaptive SD Manager continuously tunes SD parameters—choosing safe strategies to prevent OOM, maximizing throughput by learning effective strategies for each workload, and balancing speculative gains against computational overhead. In addition, to handle large batch sizes where SD offers little benefit, it employs an elastic mechanism that activates SD only when the remaining request count drops below a configurable threshold (default: 32).